# Quantum-enhanced sensing of spin-orbit coupling without fine-tuning

### v1

@@ 1 Introduction
In this paper, we show that 1D quantum wires with SOC can provide a platform for Heisenberg precision sensing over a large area in phase space. This finding provides a route for many-body quantum probe without exploiting only the vicinity of a critical point. Therefore, it opens up a way for Heisenberg precision sensing without fine-tuning. In this work we first consider the single parameter estimation scenario, where the Heisenberg scaling is shown for the cases of a single-particle probe, many-body interacting probe, and thermal probe. We then address the multi-parameter estimation case and finally present the optimal measurement basis.
@@ 2 Overview: parameter estimation
To estimate $d$ number of unknown parameters $\bm{\alpha}=(\alpha_{1},\alpha_{2},\ldots,\alpha_{d})$, the first step is to encode them in a quantum probe whose density operator $\rho_{\bm{\alpha}}$ depends on $\bm{\alpha}$. To determine these parameters, one needs to perform a measurement, which is represented by a set of projective operators $\{\Pi_{k}\}$ applied to the probe. The outcomes follow a classical probability distribution, where the probability of the $k$-th outcome is given by $p_{k}(\bm{\alpha})=\mathrm{Tr}[\Pi_{k}\rho_{\bm{\alpha}}]$. These outcomes are fed into an estimator function to infer the parameters. For this measurement setup, the uncertainty in estimating $\bm{\alpha}$ can be quantified by the covariance matrix $[\text{Cov}({\bm{\alpha}})]_{i,j}=\braket{\alpha_{i}}\braket{\alpha_{j}}-\braket{\alpha_{i}\alpha_{j}}$. Note that the diagonal elements are the single parameter variances $\sigma^{2}_{\alpha_{i}}$. The lower bound in uncertainty is set by the Craémer-Rao inequality [81, 82, 83, 84]
@@
This basis-dependent lower bound can be optimized over all possible measurements to give the ultimate in terms of the quantum Cramér-Rao bound [84]
@@
To extract scalar inequalities from the matrix inequalities, a positive weight matrix $W$ can be multiplied on both sides, followed by a trace operation. The total uncertainty is bounded by $\sum_{i}\sigma^{2}_{\alpha_{i}}\geq\text{Tr}((\mathcal{F}^{Q})^{-1})/M$, where $W$ is taken to the identity matrix. When $W$ is a zero matrix except with $W_{i,i}$, then the single parameter uncertainty can be written as $\sigma^{2}_{\alpha_{i}}\geq(\mathcal{F}^{Q})^{-1}_{i,i}/M\geq 1/M\mathcal{F}^{Q}_{i,i}$. The last term is associated with the uncertainty in $\alpha_{i}$ when all the other parameters are known. This situation is known as the single parameter estimation problem where the error in estimating one unknown parameter $\alpha$ is lower bounded by $\sigma^{2}_{\alpha}\geq 1/MF^{C}(\alpha)\geq 1/MF^{Q}(\alpha)$. Here, $F^{C}(\alpha)$ and $F^{Q}(\alpha)$ are the single parameter classical Fisher information (CFI) and quantum Fisher information (QFI), and are just the diagonal elements in CFIM and QFIM, respectively. One choice of optimal basis, in which CFI saturates the QFI, is given by the projcetors to the eigenstates of the so called symmetric logarithmic operator (SLD) $\mathcal{L}$, which are implicitly defined for a probe state $\rho$ as $\partial_{\alpha}\rho{=}(\rho\mathcal{L}{+}\mathcal{L}\rho)/2$.
@@ 3 Model: 1D quantum wire
The model considered in this work is based on the transport of electrons in a one-dimensional ballistic quantum wire in the presence of Rashba SOC [85, 86]. Low dimensional transport properties of electrons are of extreme importance in condensed matter physics, especially the emerging fields of spintronics, which explores spin degrees of freedom for information processing. SOC is quite prevalent in two-dimensional electron gas systems with strong confinement along the $z$-axis. Further confinement to a wire geometry enhances the effect of SOC [17] and has led to extensive studies on the transports of non-interacting electrons in quasi 1D systems. The effective dynamics can be described by a tight-binding lattice model, which is governed by the Hamiltonian in the following form
@@
$H_{R}=$ $-\alpha_{z}\sum_{j,\sigma,\sigma^{\prime}}c^{{\dagger}}_{j+1,\sigma}(i\sigma_{y})_{\sigma,\sigma^{\prime}}c_{j,\sigma^{\prime}}$
@@
where $\alpha_{y(z)}$ are the SOC parameters and stem from the confining potentials of the wire. The Zeeman term $H_{Z}$ accounts for the presence of an external magnetic field $\mathbf{B}=B(\sin{\theta}\cos{\phi},\sin{\theta}\sin{\phi},\cos{\theta})$,
$H_{Z}=B\sum_{j,\sigma,\sigma^{\prime}}$ $\left[(\sigma_{x})_{\sigma,\sigma^{\prime}}\sin{\theta}\cos{\phi}\right.$
$+(\sigma_{y})_{\sigma,\sigma^{\prime}}\sin{\theta}\sin{\phi}$
$\left.+(\sigma_{z})_{\sigma,\sigma^{\prime}}\cos{\theta}\right]c^{{\dagger}}_{j,\sigma}c_{j,\sigma^{\prime}}.$ (9)
In the quasi-momentum $(k)$ representation, the eigenenergies of the system is given by
$E^{(s)}(k)=(-2t\cos k+2s\,\text{sgn}(k-k_{0})\sqrt{D(k)}),$ (10)
where $s{=}{\pm}1$ indicates directions of the plane wave solutions along the wire that gives rise to two branches of energy. Here, $D(k){=}(\alpha_{y}^{2}+\alpha_{z}^{2})\sin^{2}k+B^{2}/4+B\sin k(\alpha_{y}\cos h-\alpha_{z}\sin h\sin{\phi})$, with $k_{0}{=}\sin^{-1}{(-B(\alpha_{y}\cos h-\alpha_{z}\sin h\sin{\phi})/2(\alpha_{y}^{2}+\alpha_{z}^{2}))}$.
@@ 4 Single Parameter Sensing
In low-dimensional transport experiments, accurate knowledge of the SOC parameters is desired. Therefore, we first focus on estimating the SOC parameter in the uniform Rashba field scenario along $y$- and $z$-direction, i.e. $\alpha_{y}{=}\alpha_{z}$. The non-uniform SOC case will be addressed in the next section. The goal is to analyze how well one can estimate $\alpha_{z}$ using the ground state of the system as a probe. To stay close to the experimental systems, we consider finite wires with $L$ lattice sites and study the growth of QFI with system size. As the scaling of QFI is closely connected to the gap closing of the system, we also analyze the energy gap $\Delta$ between first excited state and ground state. The external magnetic field is essential to create the gap, as the ground state is degenerate in its absence. The field can be chosen to be perpendicularly aligned to the wire, so without loss of generality, we take the wire along the $x$-direction and $\mathbf{B}$ along $z$-direction, i.e. $\theta{=}0,\phi{=}0$.
@@ 4.1 Single particle probe
Figure 1: Scaling analysis of energy gap and QFI. Probe state is the ground state of the Hamiltonian in Eqs. (6)-(9) for sensing $\alpha_{z}$ with hopping parameter $t{=}1$, $\theta{=}0$, $\phi{=}0$, and $\alpha_{y}{=}\alpha_{z}$. (a) Energy gap between first excited state and ground state $\Delta$ with $\alpha_{z}$ and $B$ for system size $L{=}100$. The four points in the parameter space are chosen as $(B,\alpha_{z})=(0.05,0.1),(0.05,0.9),(0.9,0.9),(0.9,0.05)$, denoted by the cross, diamond, plus, and circle, respectively. (b) Scaling relation of the energy gap with system size at these four parameter points, labeled by corresponding legends. In all cases numerical fit $\Delta{=}aL^{-\mu}{+}b$ shows almost quadratic scaling of the gap closing. (c) QFI of the probe with respect to $\alpha_{z}$ in the parameter space spanned by $\alpha_{z}$ and $B$ for system size $L{=}100$. The same four points are chosen to study the scaling behavior. (d) Numerical fit confirms the quadratic scaling of QFI in the whole range as $F^{Q}{\sim}L^{\beta}$ with $\beta{\approx}2$.
@@
We start the sensing analysis with the single particle ground states of the Hamiltonian in Eq.(6) as the probe. In Fig. 1, we show the behavior of the energy gap $\Delta$ and QFI $F^{Q}$ of the ground state with respect to $\alpha_{z}$. We have considered experimentally relevant values of $\alpha_{z}$ and $B$ for our analysis. Figure 1(a) shows the result for a fixed size $L=100$ where, for weak magnetic fields, the gap goes through a minima at a small value of $\alpha_{z}$. As the filed becomes stronger, the minima becomes less pronounced and the gap increases almost monotonically with $\alpha_{z}$. Interestingly, the scaling of the decline of energy gap with system size shows similar behavior across the entire range of $\alpha_{z}$ and $B$. This is shown in Fig. 1(b), where four different combinations of $\alpha_{z}$ and $B$ values in the weak and strong regimes have been chosen to analyze the scaling. In all the cases, the gap closing was found to scale almost quadratically with system size, i.e. $\Delta\sim L^{-\mu}$ with $\mu\approx 2$. Now for the QFI, Fig. 1(c) shows almost a complementary picture to Fig. 1(a), illustrating the close connection between the gap closing and the QFI. This leads to a quadratic scaling of QFI in the whole region spanned by $\alpha_{z}$ and $B$, i.e. $F^{Q}\sim L^{\beta}$ with $\beta\approx 2$. This is a very promising result from the perspective of estimating SOC parameters to great precisions. Heisenberg scaling can be achieved at the vicinity of topological phase transitions induced by infinitesimal changes in SOC parameters, but the quantum advantage is only available at criticality, which limits the the range of the application of sensors based on such criticality. This work shows that it is possible to achieve the Heisenberg limit in a wide range of the SOC parameters, owing to the favorable gap closing nature of the chosen model. We also note that for a given system size, the sensing capacity is stronger for smaller values of SOC parameters in the presence of a weak external magnetic field.
@@ 4.2 Many-body interacting probe
Figure 2: Many-body probe. Probe state is the half-filled ground state of the Hamiltonian in Eq. (11) with $t{=}1$, $\theta{=}0$, $\phi{=}0$, $B{=}0.01t$, and $\alpha_{y}{=}\alpha_{z}$. (a) Scaling of energy gap in the non-interacting case ($U{=0}$) with system size $L$ for different values of SOC parameter $\alpha_{z}$. (b) Scaling of QFI with respect to $\alpha_{z}$ in the non-interacting case for different $\alpha_{z}$ values. (c) QFI at $\alpha_{z}{=}0.1t$ as a function of interaction strength $U$. (d) QFI scaling at $\alpha_{z}{=}0.1t$ for different $U$ values.
Now we study the effects of two-body contact interactions on the sensing capability of the probe. In one-dimensional (1D) systems, such interactions are known to strongly influence the low-energy physics of many-body systems, leading to Luttinger liquid behavior. The Hamiltonian in Eq. (6) is modified to include the interaction between two fermions with different spins at the same site with strength $U$. This results in rewriting the Hamiltonian as
$H=H_{0}+H_{R}+H_{Z}+H_{U},$ (11)
@@
$H_{U}=U\sum_{j}c^{{\dagger}}_{j,\uparrow}c_{j,\uparrow}\,\,c^{{\dagger}}_{j,\downarrow}c_{j,\downarrow}.$ (12)
As a typical scenario in theoretical and experimental situations, we consider the half-filled case here, i.e. $L$ fermions on a $L$-site wire. When $U=0$, the fermionic ground state is given by the antisymmetric Slater determinant state formed by the single particle energy eigenstates {|$\{\ket{E_{n}}\}$, which can be compactly written with the symmetric group $S_{L}$ as
@@
This results in an analytical expression for the QFI [78],
@@
Top row of Fig. 2 shows the results of the many-body non-interacting probe where the external field is fixed at $B{=}0.01t$. As the single particle eigenstates are filled up to $L$-th level in the ground state, the energy gap is defined by the energy difference between the $L{+}1$ and $L$-th eigenvalues. As Fig. 2(a) shows, gap closing is still algebraic with system size $(\Delta{\sim}L^{-\mu})$, but it is less than quadratic. The exponent $\alpha$ is found to be decreasing with increasing $\alpha_{z}$. On the other hand, the QFI still scales quadratically with system size, as shown in Fig. 2(b). This shows that, the fermionic statistics in the non-interacting many-body ground state still results in the Heisenberg limit of precision for a wide range of the SOC parameter. When the interaction is non-zero, we observe that the QFI starts increasing with $U$ before attaining a saturation in the repulsive case. For attractive interaction, the QFI decreases and goes to zero. Figure 2(c) displays this behavior, while Fig. 2(d) shows that algebraic scaling of QFI is still sustained in the repulsive case. The exponential growth of the Hilbert space dimension limits our calculations in terms of system size, which can therefore be affected by finite size effects. Nevertheless, the quadratic scaling of QFI for $U{=}0$ is expectedly observed in Fig. 2(d). Surprisingly, as $U$ is increased, the decrease in exponent $\beta$ is not drastic, which indicates that the interacting probes can still provide quantum enhancement.

### v2

@@ 1 Introduction
In this paper, we show that 1D quantum wires with SOC can provide a platform for Heisenberg precision sensing over a large area in parameter space. This finding provides a route for many-body quantum probe without exploiting only the vicinity of a critical point. Therefore, it opens up a way for Heisenberg precision sensing without fine-tuning. In this work we first consider the single parameter estimation scenario, where the Heisenberg scaling is shown for the cases of a single-particle probe, many-body interacting probe, and thermal probe. We then address the multi-parameter estimation case and finally present the optimal measurement basis.
@@ 2 Overview: parameter estimation
To estimate $d$ number of unknown parameters $\bm{\alpha}=(\alpha_{1},\alpha_{2},\ldots,\alpha_{d})$, the first step is to encode them in a quantum probe whose density operator $\rho_{\bm{\alpha}}$ depends on $\bm{\alpha}$. To determine these parameters, one needs to perform a measurement, which is represented by a set of Positive Operator-Valued Measure (POVM) operators $\{\Pi_{k}\}$ applied to the probe. The outcomes follow a classical probability distribution, where the probability of the $k$-th outcome is given by $p_{k}(\bm{\alpha})=\mathrm{Tr}[\Pi_{k}\rho_{\bm{\alpha}}]$. These outcomes are fed into an estimator function to infer the parameters. For this measurement setup, the uncertainty in estimating $\bm{\alpha}$ can be quantified by the covariance matrix $[\text{Cov}({\bm{\alpha}})]_{i,j}{=}\braket{\alpha_{i}\alpha_{j}}-\braket{\alpha_{i}}\braket{\alpha_{j}}$ Note that the diagonal elements are the single parameter variances $\sigma^{2}_{\alpha_{i}}$. The lower bound in uncertainty is set by the Craémer-Rao inequality [81, 82, 83, 84]
@@
This basis-dependent lower bound can be optimized over all possible measurements to give the ultimate obtainable precision in terms of the quantum Cramér-Rao bound [84]
@@
To extract a scalar form from the matrix inequalities, a positive weight matrix $W$ can be multiplied on both sides, followed by a trace operation. In particular, for the choice of $W$ being an identity matrix one gets $\sum_{i}\sigma^{2}_{\alpha_{i}}\geq\text{Tr}((\mathcal{F}^{Q})^{-1})/M$. Another useful choice is to take a $W$ matrix with all elements being zero except $W_{i,i}{=}1$. Then the Cramér-Rao inequality becomes $\sigma^{2}_{\alpha_{i}}\geq(\mathcal{F}^{Q})^{-1}_{i,i}/M\geq 1/M\mathcal{F}^{Q}_{i,i}$, in which the last term is associated with the uncertainty in $\alpha_{i}$ when all the other parameters are known. This situation is known as the single parameter estimation problem where the error in estimating one unknown parameter $\alpha$ is lower bounded by $\sigma^{2}_{\alpha}\geq 1/MF^{C}(\alpha)\geq 1/MF^{Q}(\alpha)$. Here, $F^{C}(\alpha)$ and $F^{Q}(\alpha)$ are the single parameter classical Fisher information (CFI) and quantum Fisher information (QFI), and are just the diagonal elements in CFIM and QFIM, respectively. One choice of optimal basis, in which CFI saturates the QFI, is given by the projcetors to the eigenstates of the so called symmetric logarithmic operator (SLD) $\mathcal{L}$, which are implicitly defined for a probe state $\rho$ as $\partial_{\alpha}\rho{=}(\rho\mathcal{L}{+}\mathcal{L}\rho)/2$.
@@ 3 Model: 1D quantum wire
The model considered in this work is based on the transport of electrons in a one-dimensional ballistic quantum wire in the presence of Rashba SOC [85, 86]. Low dimensional transport properties of electrons are of extreme importance in condensed matter physics, with applications in spintronics and quantum information processing. SOC is quite prevalent in two-dimensional electron gas systems with strong confinement along the $z$-axis. Further confinement to a wire geometry enhances the effect of SOC [17] and has led to extensive studies on the transports of non-interacting electrons in quasi 1D systems. The effective dynamics can be described by a tight-binding lattice model, which is governed by the Hamiltonian in the following form
@@
$H_{R}=$ $-\alpha_{z}\sum_{j,\sigma,\sigma^{\prime}}(i\sigma_{y})_{\sigma,\sigma^{\prime}}c^{{\dagger}}_{j+1,\sigma}c_{j,\sigma^{\prime}}$
@@
where $\alpha_{y(z)}$ are the SOC parameters and stem from the confining potentials of the wire. The Zeeman term $H_{Z}$ accounts for the presence of an external magnetic field $B$ along z direction,
$H_{Z}=B\sum_{j,\sigma,\sigma^{\prime}}(\sigma_{z})_{\sigma,\sigma^{\prime}}c^{{\dagger}}_{j,\sigma}c_{j,\sigma^{\prime}}.$ (9)
@@ 4 Single Parameter Sensing
In low-dimensional transport experiments, accurate knowledge of the SOC parameters is desired. Therefore, we first focus on estimating the SOC parameter in the uniform Rashba field scenario along $y$- and $z$-direction, i.e. $\alpha_{y}{=}\alpha_{z}$. The non-uniform SOC case will be addressed in the next section. The goal is to analyze how well one can estimate $\alpha_{z}$ using the ground state of the system as a probe. To stay close to the experimental systems, we consider finite wires with $L$ lattice sites and study the growth of QFI with system size. As the scaling of QFI is closely connected to the gap closing of the system, we also analyze the energy gap $\Delta$ between first excited state and ground state. The ground state is doubly degenerate, so a Zeeman term is essential to break the degeneracy. This term need to be strong enough to capture the actual energy gap and not the Zeeman splitting. For example, one should choose $B{>}0.005t$ for $L{=}100$.
@@ 4.1 Single particle probe
Figure 1: Scaling analysis of energy gap and QFI. Probe state is the ground state of the Hamiltonian in Eqs. (6)-(9) for sensing $\alpha_{z}$ with hopping parameter $t{=}1$, and $\alpha_{y}{=}\alpha_{z}$. (a) Energy gap between first excited state and ground state $\Delta$ with $\alpha_{z}$ and $B$ for system size $L{=}100$. The four points in the parameter space are chosen as $(B,\alpha_{z})=(0.05,0.1),(0.05,0.9),(0.9,0.9),(0.9,0.05)$, denoted by the cross, diamond, plus, and circle, respectively. (b) Scaling relation of the energy gap with system size at these four parameter points, labeled by corresponding legends. In all cases numerical fit $\Delta{=}aL^{-\mu}{+}b$ shows almost quadratic scaling of the gap closing. (c) QFI of the probe with respect to $\alpha_{z}$ in the parameter space spanned by $\alpha_{z}$ and $B$ for system size $L{=}100$. The same four points are chosen to study the scaling behavior. (d) Numerical fit confirms the quadratic scaling of QFI in the whole range as $F^{Q}{\sim}L^{\beta}$ with $\beta{\approx}2$.
@@
We start the sensing analysis with the single particle ground states of the Hamiltonian in Eq.(6) as the probe. In Fig. 1, we show the behavior of the energy gap $\Delta$ and QFI $F^{Q}$ of the ground state with respect to $\alpha_{z}$. We have considered experimentally relevant values of $\alpha_{z}$ and $B$ for our analysis. Fig 1(a) shows the result for a fixed size $L=100$ where, for weak magnetic fields, the gap goes through a minima at a small value of $\alpha_{z}$. As the field becomes stronger, the minima becomes less pronounced and the gap increases almost monotonically with $\alpha_{z}$. Interestingly, the scaling of the decline of energy gap with system size shows similar behavior across the entire range of $\alpha_{z}$ and $B$. This is shown in Fig. 1(b), where four different combinations of $\alpha_{z}$ and $B$ values in the weak and strong regimes have been chosen to analyze the scaling. In all the cases, the gap closing was found to scale almost quadratically with system size, i.e. $\Delta\sim L^{-\mu}$ with $\mu\approx 2$. Now for the QFI, Fig. 1(c) shows almost a complementary picture to Fig. 1(a), illustrating the close connection between the gap closing and the QFI. This leads to a quadratic scaling of QFI in the whole region spanned by $\alpha_{z}$ and $B$, i.e. $F^{Q}\sim L^{\beta}$ with $\beta\approx 2$, see Fig. 1(d). This is a very promising result as unlike criticality-based quantum sensors, where quantum-enhanced sensitivity is only achievable at the vicinity of the phase transition, one can reach Heisenberg scaling over a wide range of parameters. This can be explained by the favorable gap closing nature of the system. We also note that for a given system size, the sensing capacity is stronger for smaller values of SOC parameters in the presence of a weak external magnetic field.
@@ 4.2 Many-body interacting probe
Now we study the effects of two-body contact interactions on the sensing capability of the probe. In one-dimensional systems, such interactions are known to strongly influence the low-energy physics of many-body systems, leading to Luttinger liquid behavior. The Hamiltonian in Eq. (6) is modified to include the interaction between two fermions with different spins at the same site with strength $U$ and between two fermions on adjacent sites with strength $V$. This results in rewriting the Hamiltonian as
$H=H_{0}+H_{R}+H_{Z}+H_{\rm int},$ (10)
@@
$H_{\rm int}=U\sum_{j}c^{{\dagger}}_{j,\uparrow}c_{j,\uparrow}+V\sum_{j,\sigma,\sigma^{\prime}}c^{{\dagger}}_{j,\sigma}c_{j,\sigma}c^{{\dagger}}_{j+1,\sigma^{\prime}}c_{j+1,\sigma^{\prime}}.$ (11)
The $U$ term accounts for an increase in the system’s energy when two fermions with opposite spins occupy the same site. For fermions in optical lattices, this term can be tuned by changing the scattering length with Feshbach resonance [87]. The $V$ term can arise due to the Coulomb repulsion, for example, between electrons confined in quantum dot arrays. As a typical scenario in theoretical and experimental situations, we consider the half-filled case here, i.e. $L$ fermions on a $L$-site wire. When $U{=}V{=}0$, the fermionic ground state is given by the antisymmetric Slater determinant state formed by the single particle energy eigenstates {|$\{\ket{E_{n}}\}$, which can be compactly written with the symmetric group $S_{L}$ as
@@
This results in an analytical QFI expression [78],
@@
Figure 2: Many-body probe. Probe state is the half-filled ground state of the Hamiltonian in Eq. (10) with $t{=}1$, $B{=}0.01t$, and $\alpha_{y}{=}\alpha_{z}$. (a) Scaling of energy gap in the non-interacting case ($U{=}V{=}0$) with system size $L$ for different values of SOC parameter $\alpha_{z}$. (b) Scaling of QFI with respect to $\alpha_{z}$ in the non-interacting case for different $\alpha_{z}$ values. (c) QFI at $\alpha_{z}{=}0.1t$ as a function of $U$ when $V{=}0$. (d) QFI scaling at $\alpha_{z}{=}0.1t$ for different $U$ values. (e) Spectrum $E$ as a function of $U$. (f) QFI at $\alpha_{z}{=}0.1t$ as a function of $V$ when $U{=}0$.
Top row of Fig. 2 shows the results of the many-body non-interacting probe where the external field is fixed at $B{=}0.01t$. As the single particle eigenstates are filled up to $L$-th level in the ground state, the energy gap is defined by the energy difference between the $L{+}1$ and $L$-th eigenvalues. As Fig. 2(a) shows, gap closing is still algebraic with system size $(\Delta{\sim}L^{-\mu})$, but it is less than quadratic. The exponent $\mu$ is found to be decreasing with increasing $\alpha_{z}$. On the other hand, the QFI still scales quadratically with system size, as shown in Fig. 2(b). This shows that, the fermionic statistics in the non-interacting many-body ground state still results in the Heisenberg limit of precision for a wide range of the SOC parameter. When only $U$ is non-zero, we observe that the QFI starts increasing with $U$ before attaining a saturation in the repulsive case. For attractive $U$, the QFI decreases and goes to zero. Fig 2(c) displays this behavior, while Fig. 2(d) shows that algebraic scaling of QFI is still sustained in the repulsive case. The exponential growth of the Hilbert space dimension limits our calculations in terms of system size, which can therefore be affected by finite size effects. Nevertheless, the quadratic scaling of QFI for $U{=}0$ is expectedly observed in Fig. 2(d). Surprisingly, as $U$ is increased, the decrease in exponent $\beta$ is not drastic, which indicates that the interacting probes can still provide quantum enhancement. This can be attributed to the fact that the presence of $U$ in Eq. (11) does not open up the ground state energy gap as shown in the band structure in Fig. 2(e). The fermions in the ground state stay localized at different sites. In the other case, when $V$ is non-zero and $U{=}0$, the QFI steadily decreases (see Fig. 2(f)). We numerically observe that quadratic scaling of QFI survives for small values of $V$ and disappears near $V{\approx}t$.